# Probing quantum chaos through singular-value correlations in sparse non-Hermitian SYK model

### v1

@@ Abstract
Figure 2: (Top) Variation of $\braket{r_{\upsigma}}$-value with various level of sparsity $p$. The size of the circle varies with $N$ i.e., the largest $N$ corresponds to the leftmost plot for each color. At low sparsity, the $\braket{r_{\upsigma}}$ matches with the corresponding random matrix ensemble until a threshold $p_{\mathrm{crit}}$ is reached, after which $\braket{r_{\upsigma}}$ drops to the Poissonian value, indicating a hint of integrability. The estimated errors for the values obtained is smaller than the markers in the plot. (Bottom) The finite-size dependence of the threshold or the critical value $p_{\mathrm{crit}}$. It decreases as the system size increases. The decrease is $1/N^{3}$ with a large system size, as given in (6).
Spectral and Singular form factor: In the realm of quantum chaos, the Spectral Form Factor (SFF) serves as a pivotal tool for probing the full spectrum of eigenvalue correlations, encompassing both short and long-range interactions. The SFF is defined as the Fourier transform of two-point eigenvalue correlation pairs Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997)
$\mathrm{SFF}(t)=\frac{1}{L^{2}}\braket{}{Z_{E}(it)}{{}^{2}}=\frac{1}{L^{2}}\bigg{\langle}\Big{|}\sum_{n}e^{-iE_{n}t}\Big{|}^{2}\bigg{\rangle}\,,$ (7)
where $Z_{E}(it)=\sum_{n}e^{-iE_{n}t}$ is the infinite-temperature partition function for the energy eigenvalues, and $L$ is the dimension of the Hilbert space. Here the angle brackets indicate an ensemble average, which is particularly relevant in the context of a disordered system since SFF is not self-averaging Prange (1997). Throughout our discussion, we consider the quenched disorder, instead of annealed one, which is less physical in this scenario. Interestingly, in the context of a disorder-free SYK model Lau et al. (2021), the SFF reveals indications of integrability, marking a departure from the chaotic behavior Ozaki and Katsura (2024).
For systems exhibiting quantum chaos, the SFF displays a characteristic dip-ramp-plateau pattern. The initial dip, evident in the early-time regime, is a non-universal trait influenced by the detailed nature of the energy spectrum Agam et al. (1995). This dip extends until a timescale known as the Thouless time ($t_{\mathrm{Th}}$) Gharibyan et al. (2018); Nosaka et al. (2018) or ramp time Cotler et al. (2017) beyond which the energy difference surpasses the inverse of ($t_{\mathrm{Th}}$), referred to as the Thouless energy Thouless (1977). Subsequently, a ramp emerges at $t_{\mathrm{Th}}$, indicative of eigenvalue repulsion, a phenomenon that mirrors the SFF behavior predicted by RMT Brézin and Hikami (1997); Cotler et al. (2017) and underscores the chaotic dynamics of the Hamiltonian. Moreover, this ramp hints at the presence of a holographic dual to the corresponding theory. Beyond the Heisenberg time Erdős and Knowles (2015); Prange (1997), which scales inversely to the mean level spacing, the SFF reaches a plateau that depends on the system size. In contrast, integrable systems typically bypass the ramp stage and saturate directly from the dip to the plateau regime. SFF in sparse SYK models has been thoroughly investigated in the literature García-García et al. (2021); Tezuka et al. (2023); Anegawa et al. (2023); Orman et al. (2024).
@@
where $Z_{\upsigma}(it)=\sum_{n}e^{-i\sigma_{n}t}$ is the infinite-temperature partition function of singular values, which can be further expressed as the overlap between the time-evolved right or left singular vectors governed by the effective Hamiltonian $\sqrt{H^{\dagger}H}$ or $\sqrt{HH^{\dagger}}$ respectively Roccati et al. (2024).
Figure 3: Behavior singular form factor $\upsigma$FF$(t)$ for various sparsity parameter $p$. The linear ramp ceases to exist as the sparsity increases. The system parameters are $N=26$ (GUE) with $1000$ Hamiltonian realizations. Here we choose the Gaussian filter (10) with $\alpha=3.27$, which lies within the prescribed range in Orman et al. (2024).
Figure 3 shows the distinct dip-ramp-plateau pattern exhibited by $\upsigma$FF for the non-Hermitian model (2), across varying degrees of sparsity. In this study, we consider a system size of $N=26$ and analyze the outcomes from $1000$ independent realizations of the Hamiltonian to ensure robust statistical analysis. To reduce the oscillations in the dip region, we use the filter function $Y_{\upsigma}(\alpha,t)$ Gharibyan et al. (2018) such that
@@
Here the filter function for the singular values is chosen to be Gaussian, which we refer to as the Gaussian singular filter function. It is defined as
$|Y_{\upsigma}(\alpha,t)|^{2}=\bigg{|}\sum_{n}e^{-\alpha\sigma_{n}^{2}}\,e^{-i\sigma_{n}t}\bigg{|}^{2}\,.$ (10)
As shown for Hermitian systems, given the energy spectrum is centered around $E=0$, such a filter function modifies the density of states $\rho(E)\rightarrow e^{-\alpha E^{2}}\rho(E)$ such that the $\upsigma$FF receives the dominant contribution from the bulk of the energy spectrum Gharibyan et al. (2018); Nosaka et al. (2018). Here $\alpha$ is a free parameter and controls the Gaussianity of the spectrum. This filtering avoids the unfolding which can otherwise be studied using the connected unfolded SFF Nosaka et al. (2018). For our computation, we employ the filter function on the singular values by choosing $\alpha=3.27$, which lies inside the range given in Orman et al. (2024) (dependent on $N$). Filter functions for SFF have also recently been shown to be associated with quantum channels Matsoukas-Roubeas et al. (2023b).
A notable observation is a parallelism between the behavior of $\upsigma$FF in non-Hermitian systems and the SFF in Hermitian systems. This similarity underscores the efficacy of $\upsigma$FF as a robust indicator for quantum chaos in non-Hermitian frameworks. In regimes of lower sparsity, the presence of the ramp is clearly evident, signifying the retention of chaotic dynamics within this parameter space. It remains pronounced until a critical sparsity threshold is reached. Beyond this juncture, the linear ramp—a signature trait of non-integrable systems—ceases to manifest, indicating a shift towards regularity and a departure from chaotic behavior.
Figure 4 shows the variation of Thouless time ($t_{\mathrm{Th}}$), for $N=26$, with sparseness parameter $p$. The $t_{\mathrm{Th}}$ is calculated by considering the time at which the $\upsigma\mathrm{FF}$ of the system considered first intersects the SFF of the GUE random matrices Cotler et al. (2017). To determine the intersection point precisely we consider the the following fractional error function Gharibyan et al. (2018)
$\epsilon(t):=\bigg{|}\frac{\upsigma\mathrm{FF}(t)-\upsigma_{\mathrm{ramp}}(t)}{\upsigma_{\mathrm{ramp}}(t)}\bigg{|}\,,$ (11)
indicating the deviation of the computed $\upsigma\mathrm{FF}$ to the linear-fitted function for the ramp $\upsigma_{\mathrm{ramp}}(t)$ for $p=1$ (dense SYK model (1)). We consider the value $t$ as the Thouless time or ramp time $t_{\mathrm{Th}}$, where the fractional error $\epsilon(t)$ reaches $20\%$. The red line shows the fitting $t_{\mathrm{Th}}\approx a/p^{b}+c$ with $a=1.923$, $b=0.8767$ and $c=0.007$. Curiously, this fitting follows closely to the $1/p$ scaling in contrast to the $1/p^{2}$ scaling in Hermitian systems Orman et al. (2024). Such distinction is possible since the singular values and the eigenvalues are inherently different. Yet, we do not rule out the possibility of the scaling of the exponent $b$ with the system size.
@@
In Hermitian systems, a key observation is that the singular values correspond to the absolute values of the eigenvalues. This distinction is crucial, as it implies that the SFF and $\upsigma$FF, exhibit fundamentally different behaviors. This difference is illustrated in the inset of Fig. 4, which depicts the dynamics of SFF and $\upsigma$FF for the dense Hermitian SYK4 model. The model is derived by setting $M_{abcd}=0$ for all $\{a,b,c,d\}$ in its non-Hermitian counterpart (Eq. (1)). Replacing $\sigma_{n}$ by eigenvalues $E_{n}$ in (9), we get the analogous definition of filtered SFF.
@@
Table 2: The Table shows the comparison between the singular value $\braket{r_{\upsigma}}$-value and the $\braket{r}$-values for the dense Hermitian SYK model for different system sizes $N=20$ (GSE), $N=22$ (GUE), $N=24$ (GOE). We take 1000 Hamiltonian realizations for each case. The Hermitian model is obtained by setting $M_{abcd}=0$ for all $\{a,b,c,d\}$ in (1). Here the $\braket{r_{+}}$-value includes the positive eigenvalues while $\braket{r_{-}}$-value includes the negative eigenvalues only. They are to be matched with the $\braket{r}_{\mathrm{RMT}}$ values in Table 1 and Atas et al. (2013). The $\braket{r_{\upsigma}}$-values are compared with the $\braket{r}$-value of $2$ blocks (marked in blue) Orman et al. (2024). Explanation is given in the text.
@@
Intriguingly, the Hermitian system exhibits a unique characteristic in its $\braket{r_{\upsigma}}$-value as well. As depicted in Table 2, there is a notable distinction when comparing the $\braket{r_{\upsigma}}$-value to the $\braket{r}$-values within the identical Hermitian SYK4 model. Additionally, the $\braket{r_{+}}$-value and $\braket{r_{-}}$-value were computed separately, considering only the positive and negative eigenvalues of the Hamiltonian, respectively. These ratios align with the overall $\braket{r}$-value, which is anticipated since it solely accounts for the spacing between consecutive eigenvalues. Nonetheless, these individual ratios as well as the overall $\braket{r}$-value do not correspond to the $\braket{r_{\upsigma}}$-value, which is identified as the $\braket{r}$-value for two separate blocks before symmetry resolution. This discrepancy arises because, in Hermitian systems, the singular values are determined by the absolute magnitudes of the eigenvalues, leading to the computation of the $\braket{r_{\upsigma}}$-value mirroring the $\braket{r}$-value for bifurcated blocks Giraud et al. (2022); Kawabata et al. (2023); Orman et al. (2024). We anticipate such bifurcation has a similar effect of having the presence of conserved charges Iyoda et al. (2018); Sá and García-García (2022). Curiously, this phenomenon is exclusive to Hermitian and anti-Hermitian (with $J_{abcd}=0$, $M_{abcd}\neq 0$) systems and echoes the discrepancy between the $\upsigma$FF and SFF as shown in the inset of Fig. 4.
It is imperative to differentiate the behavior of $\upsigma$FF from the dissipative spectral form factor (DSFF), especially since DSFF has been instrumental in deciphering chaos in non-Hermitian and open systems Li et al. (2021); Cipolloni and Grometto (2024); Li et al. (2024). The $\upsigma$FF is characterized by its reliance on real singular values, whereas the DSFF encompasses the complex eigenspectrum, establishing a fundamental difference between the two. The DSFF reveals a unique quadratic ramp-plateau structure in non-Hermitian systems Li et al. (2021, 2024). Meanwhile, the $\upsigma$FF, while also mirroring the SFF exactly for Hermitian systems upon the appropriate identification of the singular vectors, preserves the linear ramp-plateau configuration. It, thus, serves as an effective marker for the transition between integrable and chaotic phases. Therefore, despite their inherent differences, both metrics are apt for exploring the nuances of non-Hermitian and open quantum systems.
@@
Figure 5: Behaviour of the singular complexity with time for dense ($p=1$) non-Hermitian SYK model (1). The system parameter is $N=26$ with $1000$ random Hamiltonian realizations taken. Inset shows the early time behavior of the complexity, which increases quadratically followed by linear growth and saturation.
Figure 6: Behaviour of the saturation of the singular complexity with different sparseness for $N=26$ with $1000$ random Hamiltonian realizations.
Spectral complexity for singular values - Singular complexity: A related quantity, capturing the long-range spectral correlation is motivated by considering the spectral complexity Iliesiu et al. (2022). Within the framework of the AdS/CFT duality Maldacena (1998); Witten (1998), spectral complexity finds its counterpart in the volumetric measure of the Einstein-Rosen bridge, serving as an analog for the computational and circuit complexity Jefferson and Myers (2017). In a similar spirit, we define singular complexity at finite temperature $T=1/\beta$ as
$C_{\upsigma}(t)=\frac{1}{Z_{2\upsigma}(\beta)L}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}e^{-\beta(\sigma_{i}+\sigma_{j})}\,,$ (13)
@@
designed to be suitable for non-Hermitian systems. Here $Z_{\upsigma}(\beta)=\sum_{n}e^{-\beta\sigma_{n}}$ is the thermal singular partition function, constructed from the singular values of the corresponding non-Hermitian Hamiltonian. Assuming no degeneracy in the singular value spectrum, the derivative of the singular complexity is related to the $\upsigma$FF as
@@
in a similar spirit to spectral complexity Iliesiu et al. (2022); Erdmenger et al. (2023) which can be checked directly using (8). Focusing on the infinite temperature $\beta=0$ limit, (13) takes a simpler form

### v2

@@ Abstract
Figure 2: (Top) Variation of $\braket{r_{\upsigma}}$-value with various level of sparsity $p$. The size of the circle varies with $N$ i.e., the largest $N$ corresponds to the leftmost plot for each color. At low sparsity, the $\braket{r_{\upsigma}}$ matches with the corresponding random matrix ensemble until a threshold $p_{\mathrm{crit}}$ is reached, after which $\braket{r_{\upsigma}}$ drops to the Poissonian value, indicating a hint of integrability. The estimated errors for the values obtained are smaller than the markers in the plot. (Bottom) The finite-size dependence of the threshold or the critical value $p_{\mathrm{crit}}$. It decreases as the system size increases. The decrease scales as $1/N^{3}$ with large system size, as given in (4).
Spectral and Singular form factor: Contrary to the singular value statistics, the SFF probes the long-range eigenvalue correlations. It is defined as the Fourier transform of two-point eigenvalue correlation pairs Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997). For systems exhibiting quantum chaos, the SFF displays a characteristic dip-ramp-plateau pattern. The initial dip, evident in the early-time regime, is a non-universal trait influenced by the detailed nature of the energy spectrum Agam et al. (1995). This dip extends until a timescale known as the Thouless time ($t_{\mathrm{Th}}$) Gharibyan et al. (2018); Nosaka et al. (2018) or ramp time Cotler et al. (2017) beyond which the energy difference surpasses the inverse of $t_{\mathrm{Th}}$, referred to as the Thouless energy Thouless (1977). Subsequently, a ramp emerges at $t_{\mathrm{Th}}$, indicative of eigenvalue repulsion, a phenomenon that mirrors the SFF behavior predicted by RMT Brézin and Hikami (1997); Cotler et al. (2017) and underscores the chaotic dynamics of the Hamiltonian. Moreover, this ramp hints at the presence of a holographic dual to the corresponding theory. Beyond the Heisenberg time Erdős and Knowles (2015); Prange (1997), which scales inversely to the mean level spacing, the SFF reaches a plateau that depends on the system size. In contrast, integrable systems typically bypass the ramp stage and saturate directly from the dip to the plateau regime. SFF in sparse SYK models has been thoroughly investigated in recent literature García-García et al. (2021); Tezuka et al. (2023); Anegawa et al. (2023); Orman et al. (2024). Several other generalizations for SFF in non-Hermitian and open systems have also been proposed García-García et al. (2023a); Xu et al. (2021); Zhou et al. (2024).
@@
where $Z_{\upsigma}(it)=\sum_{n}e^{-i\sigma_{n}t}$ is the infinite-temperature partition function of singular values, which can be further expressed as the overlap between the time-evolved right or left singular vectors governed by the effective Hamiltonian $\sqrt{H^{\dagger}H}$ or $\sqrt{HH^{\dagger}}$ respectively Roccati et al. (2024). Here the angle brackets indicate an ensemble average, which is particularly relevant in the context of a disordered system since $\upsigma$FF is not self-averaging Prange (1997).
Figure 3: Behavior of $\upsigma$FF$(t)$ for various sparsity parameter $p$. The linear ramp ceases to exist as the sparsity increases. The system parameters are $N=26$ (GUE) with $1000$ Hamiltonian realizations. Here we choose the Gaussian filter with $\alpha=3.27$, which lies within the prescribed range in Orman et al. (2024).
Figure 3 shows the distinct dip-ramp-plateau pattern exhibited by $\upsigma$FF for the Hamiltonian (1), across varying degrees of sparsity. In this work, we consider a system size of $N=26$ and analyze the outcomes from $1000$ independent realizations of the Hamiltonian to ensure robust statistical analysis. To reduce the oscillations in the dip region, we use the filter function $Y_{\upsigma}(\alpha,t)$ Gharibyan et al. (2018) such that
@@
Here the Gaussian filter function for the singular values is defined as $|Y_{\upsigma}(\alpha,t)|^{2}=|\sum_{n}e^{-\alpha\sigma_{n}^{2}}\,e^{-i\sigma_{n}t}|^{2}$. Similar to the eigenspectrum Gharibyan et al. (2018); Nosaka et al. (2018), it ensures that the dominant contribution is received from the bulk of the singular value spectrum. Here $\alpha$ is an $N$-dependent parameter that controls the Gaussianity of the spectrum. This filtering is performed for the usual spectrum, without unfolding it. For the latter case, an equivalent of connected unfolded SFF García-García et al. (2018a); Nosaka et al. (2018) can be defined. We employ the filter function on the singular values by choosing $\alpha=3.27$, which lies inside the range given in Orman et al. (2024) for $N=26$. Filter functions for SFF have also recently been shown to be associated with quantum channels Matsoukas-Roubeas et al. (2023b).
A notable observation in Fig. 3 is a parallelism between the behavior of $\upsigma$FF in non-Hermitian systems and the SFF in Hermitian systems. This similarity underscores the efficacy of $\upsigma$FF as a robust indicator for quantum chaos in non-Hermitian frameworks. In regimes of lower sparsity, the presence of the ramp is evident, signifying the retention of chaotic dynamics within this parameter space. It remains pronounced until a critical sparsity threshold is reached. Beyond this juncture, the linear ramp—a signature trait of non-integrable systems—ceases to manifest, indicating a shift towards integrability and a departure from chaotic behavior.
Figure 4 shows the variation of Thouless time ($t_{\mathrm{Th}}$), for $N=26$, with sparseness parameter $p$. The $t_{\mathrm{Th}}$ is calculated by considering the time at which the $\upsigma\mathrm{FF}$ of the system considered first intersects the SFF of the GUE random matrices Cotler et al. (2017). To determine the intersection point precisely, we consider the fractional error function $\epsilon(t):=\Big{|}\frac{\upsigma\mathrm{FF}(t)-\upsigma_{\mathrm{ramp}}(t)}{\upsigma_{\mathrm{ramp}}(t)}\Big{|}$ Gharibyan et al. (2018) indicating the deviation of the computed $\upsigma\mathrm{FF}$ to the linear-fitted function for the ramp $\upsigma_{\mathrm{ramp}}(t)$ for $p=1$. We consider the value $t$ as the Thouless time or ramp time $t_{\mathrm{Th}}$, where the fractional error $\epsilon(t)$ reaches $20\%$. The red line shows the fitting $t_{\mathrm{Th}}\approx a/p^{b}+c$ with $a=1.923$, $b=0.8767$ and $c=0.007$. Curiously, this fitting follows closely to the $1/p$ scaling in contrast to the $1/p^{2}$ scaling in Hermitian systems Orman et al. (2024). Such distinction is possible since the singular values and the eigenvalues are inherently different (see the Supplemental Material sup S1). Yet, we do not rule out the possibility of the scaling of the exponent $b$ with $N$.
@@
Spectral complexity for singular values - Singular complexity: Within the framework of the AdS/CFT correspondence Maldacena (1998); Witten (1998), a particular spectral measure has been proposed to capture the essence of long-range spectral correlations. This measure, known as spectral complexity, is proposed to be a dual quantity of the Einstein-Rosen bridge Iliesiu et al. (2022) in the gravitational theory, serving as an analog for the computational and circuit complexity Jefferson and Myers (2017). In a similar spirit, we define singular complexity at infinite temperature as
$C_{\upsigma}(t)=\frac{1}{L^{2}}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}\,,$ (7)
designed to be suitable for non-Hermitian systems. At early times, it grows quadratically, followed by a linear growth and plateau regime (see Supplemental Material sup S2 for more details). The early time behavior mirrors the behavior of spread complexity Balasubramanian et al. (2022); Erdmenger et al. (2023) based on the Krylov space approach Parker et al. (2019). The plateau value can be obtained analytically by taking the long-time average
[…]
which only depends on the difference between the singular values. In the Hermitian case, such saturation values have been studied for quantum billiards Camargo et al. (2024a) and mixed-field Ising model Camargo et al. (2024b), capturing the spectral rigidity. Integrable systems saturate at higher values compared to chaotic ones due to the presence of level repulsion, as evident from the equivalent eigenspectrum of (8).
Figure 5: Behavior of the saturation of the singular complexity with different sparseness for $N=26$ with $1000$ random Hamiltonian realizations.
Figure 5 depicts the saturation of singular complexity at varying levels of sparsity, which maintains a magnitude order of $O(10^{4})$ until it encounters a critical threshold. Beyond this threshold, there is a marked escalation at lower sparsity levels signaling a level-repulsion between the singular values. This critical value is dependent on the system size $N$, and for $N=26$, it occurs within the sparsity regime $0.003\lesssim p_{\mathrm{crit}}\lesssim 0.004$. This value resembles the expected $\braket{r_{\upsigma}}$-value shown in Fig. 2.
Exhibiting chaotic dynamics is considered to be a crucial constraint that possesses a holographic dual geometry. While this discussion does not explore the specifics of such geometry, some proposals exist in the literature García-García and Godet (2021); García-García et al. (2023b); Cai et al. (2022); Cao and Ge (2024). The breakdown of chaotic nature suggests that such dual geometry corresponding to a non-Hermitian system may be compromised at low sparsity.
Conclusion and Outlook: Drawing on the foundational work Kawabata et al. (2023) on the applicability of SVD in non-Hermitian systems, our study addresses the quantum chaotic attributes of the sparse SYK model across varying levels of sparsity. The complex eigenvalues inherent to non-Hermitian Hamiltonians pose challenges to the generalization of traditional eigenvalue statistics. This complexity persists despite strides made with complex spacing ratios Sá et al. (2020) and the DSFF Li et al. (2021, 2024). Nonetheless, recent advancements propose a novel perspective by focusing on singular values as opposed to eigenvalues within the non-Hermitian framework. Unlike their complex counterparts, singular values are inherently real and non-negative, aligning with the real eigenvalues in Hermitian systems upon appropriate alignment with singular vectors, as detailed in the Supplemental Material sup S1.
Our investigation underscores that the statistical analysis of singular values, which elucidates short-range correlations, coupled with the $\upsigma$FF and singular complexity, illuminating long-range correlations among singular values, serve as potent indicators of the chaos within non-Hermitian systems. As such, exploring the holographic bulk dual of singular complexity may offer insights into a potential holographic counterpart for non-Hermitian systems, especially the effective range of $k$ in (4) in relation to $2d$ de Sitter (dS2) gravity García-García et al. (2023b), which we leave for future work. Moreover, these tools adeptly reflect the shift towards integrability as sparsity intensifies. Such transition differs from other chaotic-integrable transitions by two-point interactions observed in the SYK model García-García et al. (2018b). We envision that these features will be exhibited in other correlation measures such as higher spacing ratios Srivastava et al. (2018); Tekur et al. (2018); Shir et al. (2023) and the number variance García-García et al. (2023a); Gharibyan et al. (2018); Evers and Mirlin (2008); García-García and Verbaarschot (2016).
Parallel research avenues have explored the efficacy of sparse Hamiltonians Xu et al. (2020); Tezuka et al. (2023) within quantum processors Jafferis et al. (2022); Kobrin et al. (2023), maintaining chaotic dynamics to a substantial degree of sparsity. Notably, the SYK model retains its holographic duality even at minimal sparsity levels García-García et al. (2021); Orman et al. (2024). Our study contributes to the burgeoning field of non-Hermitian systems in the realm of open-system dynamics, proposing a framework for the extrapolation of quantum chaotic properties within these systems. This might serve as the benchmark results for the simulation of sparse non-Hermitian systems on quantum processors.
Acknowledgements: We would like to thank Hugo A. Camargo, Adolfo del Campo, Aurélia Chenu, Antonio M. García-García, Luca V. Iliesiu, Yiyang Jia, Hosho Katsura, and Tadashi Takayanagi for fruitful discussions, comments, and suggestions on the draft. Numerical computations were performed in the Sushiki workstation using the computational facilities of YITP. P.N. thanks the Berkeley Center for Theoretical Physics (BCTP), University of California for hosting him through the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program of Japan Science and Technology Agency (JST), Grant No. JPMJAP2318 during the final stages of the work. This work is supported by the Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05190 (P.N.) and JP21H05182, JP21H05185 (M.T.). The Yukawa Research Fellowship of T.P. is supported by the Yukawa Memorial Foundation and JST CREST (Grant No. JPMJCR19T2). The work of M.T. was partially supported by the JSPS Grants-in-Aid for Scientific Research (KAKENHI) Grants No. JP20K03787 and JST CREST (Grant No. JPMJCR24I2).
Authors’ names are listed in alphabetical order.
Supplemental Materials:
Probing quantum chaos through singular-value correlations
in sparse non-Hermitian SYK model
Pratik Nandy, Tanay Pathak, and Masaki Tezuka
Appendix A S1. From $\upsigma$FF to SFF: Reduction to Hermitian systems
@@
In Hermitian systems, a key observation is that the singular values correspond to the absolute values of the eigenvalues. This distinction is crucial, as it implies that the SFF and $\upsigma$FF, exhibit fundamentally different behaviors. This difference is illustrated in the inset of Fig. 4, which depicts the dynamics of SFF and $\upsigma$FF for the dense Hermitian SYK4 model. The model is derived by setting $x_{abcd}=1$ and $M_{abcd}=0$ for all $\{a,b,c,d\}$ in its non-Hermitian counterpart (Eq. (1)). Replacing $\sigma_{n}$ by eigenvalues $E_{n}$ in Eq. (6), we get the analogous definition of filtered SFF.
@@
Table 2: The Table shows the comparison between the singular value $\braket{r_{\upsigma}}$-value and the $\braket{r}$-values for the dense Hermitian SYK model for different system sizes $N=20$ (GSE), $N=22$ (GUE), $N=24$ (GOE). We take 1000 Hamiltonian realizations for each case. The Hermitian model is obtained by setting $x_{abcd}=1$ and $M_{abcd}=0$ for all $\{a,b,c,d\}$ in the Hamiltonian (1). Here the $\braket{r_{+}}$-value includes the positive eigenvalues while $\braket{r_{-}}$-value includes the negative eigenvalues only. They are to be matched with the $\braket{r}_{\mathrm{RMT}}$ values in Table 1 and Atas et al. (2013). The $\braket{r_{\upsigma}}$-values are compared with the $\braket{r}$-value of $2$ blocks (marked in blue) Orman et al. (2024). Explanation is given in the text.
@@
Intriguingly, the Hermitian system exhibits a unique characteristic in its $\braket{r_{\upsigma}}$-value. As depicted in Table 2, there is a notable distinction when comparing the $\braket{r_{\upsigma}}$-value to the $\braket{r}$-values within the identical Hermitian SYK4 model. Additionally, the $\braket{r_{+}}$-value and $\braket{r_{-}}$-value were computed separately, considering only the positive and negative eigenvalues of the Hamiltonian, respectively. These ratios align with the overall $\braket{r}$-value, which is anticipated since it solely accounts for the spacing between consecutive eigenvalues. Nonetheless, these individual ratios as well as the overall $\braket{r}$-value do not correspond to the $\braket{r_{\upsigma}}$-value, which is identified as the $\braket{r}$-value for two separate blocks before symmetry resolution. This discrepancy arises because, in Hermitian systems, the singular values are determined by the absolute magnitudes of the eigenvalues, leading to the computation of the $\braket{r_{\upsigma}}$-value mirroring the $\braket{r}$-value for bifurcated blocks Giraud et al. (2022); Kawabata et al. (2023); Orman et al. (2024). We anticipate such bifurcation has a similar effect of having the presence of conserved charges Iyoda et al. (2018); Sá and García-García (2022). Curiously, this phenomenon is exclusive to Hermitian and anti-Hermitian (with $J_{abcd}=0$, $M_{abcd}\neq 0$) systems and echoes the discrepancy between the $\upsigma$FF and SFF as shown in the inset of Fig. 4.
Appendix B S2. Behavior of singular complexity
@@
Figure 6: Behaviour of the singular complexity with time for dense ($p=1$) non-Hermitian SYK model. The system parameter is $N=26$ with $1000$ random Hamiltonian realizations taken. Inset shows the early time behavior of the complexity, which increases quadratically followed by linear growth and saturation.
Here we present the detailed behavior of singular complexity Iliesiu et al. (2022). It is defined at finite temperature $T=1/\beta$ as
$C_{\upsigma}(t)=\frac{1}{Z_{2\upsigma}(\beta)L}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}e^{-\beta(\sigma_{i}+\sigma_{j})}\,.$ (10)
@@
Here $Z_{\upsigma}(\beta)=\sum_{n}e^{-\beta\sigma_{n}}$ is the thermal singular partition function, constructed from the singular values of the corresponding non-Hermitian Hamiltonian. In the case of degeneracies, we keep the degenerate spectrum. The derivative of the singular complexity is related to the $\upsigma$FF as
@@
in a similar spirit to spectral complexity Iliesiu et al. (2022); Erdmenger et al. (2023) which can be checked. Focusing on the infinite temperature $\beta=0$ limit, (10) takes a simpler form as in the main text: